# JSJ decompositions of knot exteriors, Dehn surgery and the $L$-space conjecture

### v3

@@ 1. Introduction
We call a closed, connected, orientable, irreducible $3$-manifold $LO$ if its fundamental group is left-orderable, $CTF$ if it admits a co-oriented taut foliation, and $NLS$ if it is not a Heegaard Floer $L$-space. The $L$-space conjecture contends that these three conditions are equivalent ([7], [36]).
@@
Since cable knots are prime the following corollary is immediate. See Figure 1.
@@
All rational surgeries on a composite knot are $LO$ and $NLS$. ∎
@@
Suppose that $K$ is a satellite $L$-space knot. Then the JSJ graph of $K$ is a rooted interval. ∎
@@
Suppose that $K$ is a satellite knot which admits an irreducible rational surgery that is not $LO$. Then the JSJ graph of $K$ is a rooted interval. ∎
@@
We note that the hypothesis in Theorem 3.13 that $K(r)$ is irreducible is necessary, though not particularly restrictive. A fundamental result of Scharlemann states that surgery on a satellite knot $K$ yields a reducible manifold if and only if $K$ is a cable knot and the surgery slope is the cabling slope ([52]; also see §2.4). As such a surgery has a lens space summand, its fundamental group is not left-orderable.
@@
The key result that is used to prove Theorems 3.1 and 3.10 is the fact that the meridional slope of a non-trivial knot $K$ is both $LO$-detected and $NLS$-detected [8] (see Theorem 2.4). (We refer the reader to §2.1 for the definitions of a slope on $\partial X(K)$ being $LO$-, $NLS$- and $CTF$-detected.) The proof of Theorem 3.3 uses the more general fact that any slope on $\partial X(K)$ at distance 1 from the longitude, i.e. any slope of the form $1/n,n\in{\mathbb{Z}}$, is $LO$- and $NLS$-detected. In [8, Conjecture 1.6] we conjectured that the latter also holds for $CTF$-detection. In particular,
@@
In [8], we showed that Conjecture 1.1 holds for fibred knots. This allows us to partially extend our results on $LO$ and $NLS$ surgeries to $CTF$ surgeries in the case that the companion knots are fibred. See Theorems 4.1 and 4.2 for the precise statements; these are analogues of Theorems 3.1 and 3.10 respectively. In particular, we recover the following result of Delman and Roberts, which is an immediate consequence of [13, Theorem 6.1].
@@
In addition to what we have mentioned above, it is known that if a satellite knot $K=P(K_{0})$ is an $L$-space knot, then both $K_{0}$ and $P(U)$ are $L$-space knots [27]. One therefore expects the $LO$ analogue of this to hold also. In this direction, we prove the following.
@@
Suppose that $K$ is a satellite knot with a persistently foliar companion whose associated pattern has winding number $w\geq 1$. Then $K(r)$ is $CTF$ for any $r\in\mathbb{Q}$ which is not the cabling slope if $K$ is a cable knot. ∎
@@
If $K$ is a positive satellite $L$-space knot with pattern of winding number $w$ then $K(r)$ is $CTF$ for each rational $r\in(-\infty,w^{2})$. ∎
@@
At most two rational surgeries on a non-trivial knot yield integer homology $3$-spheres that are not $LO$, and the associated slopes are contained in either $\{-1,1\}$, $\{-1,-1/2\}$, or $\{1/2,1\}$. ∎
@@ Acknowledgements.
The authors wish to thank Duncan McCoy and Patricia Sorya for helpful remarks related to Theorem 3.10.
@@ 2.1. Slope detection and gluing results
  • $NLS$-detected (in $M$) if it lies in the closure of the set of slopes whose associated Dehn filling is not an $L$-space ([48, §7.2]).
@@
It is is not known whether a Dehn filling $M(\alpha)$ of a knot manifold $M$ being $CTF$ implies that the slope $\alpha$ is $CTF$-detected, though this is thought to be true (and is true if $M$ is a Seifert fibre space [3]). If so, the condition that $\alpha$ is $CTF$-detected will be strictly weaker than the condition that $M(\alpha)$ be $CTF$: The meridional slope $\mu$ of a non-trivial fibred knot is $CTF$-detected (cf. Theorem 2.4) though the $\mu$-Dehn filling is $S^{3}$ and therefore not $CTF$.
@@
The longitudinal slope of a knot manifold is normal-$\ast$-detected, where *** denotes either $NLS$, $CTF$, or $LO$.
@@
Suppose that $W=M_{1}\cup_{f}M_{2}$ where $M_{1},M_{2}$ are knot manifolds and $f:\partial M_{1}\xrightarrow{\;\cong\;}\partial M_{2}$. If $f$ identifies a normal-$\ast$-detected slope on the boundary of $M_{1}$ with a normal-$\ast$-detected slope on the boundary of $M_{2}$, then $W$ has property normal-$\ast$.
@@ 2.2. Conventions on satellite knots
For a coprime pair $p$ and $q\geq 0$, define $P(p/q)$ to be the $p\mu+q\lambda$ Dehn filling of $M$ along $\partial X(K)$. Equivalently, $M(p/q)$ is the $p/q$-Dehn surgery on the knot $P$ in $V$. We often denote $M(p/q)$ by $P(p/q)$. Similarly, we use $K(p/q)$ to denote the $p\mu+q\lambda$ Dehn filling of $X(K)$ along $\partial X(K)$. Then
@@
Let $K=P(K_{0})$ be a satellite knot with pattern $P$ and companion $K_{0}$. Fix $r\in\mathbb{Q}$. If $T=\partial X_{0}$ is incompressible in $K(r)$ and $\lambda_{P(r)}$ is ***-detected in $X_{0}$, then $K(r)$ has property ***, where *** is either $LO,NLS$, or $CTF$.
@@ 2.3. Conventions on cable knots
Suppose $K=P(K_{1})$. This gives a decomposition $X(K)=M_{1}\cup_{T_{1}}X_{1}$, where $M_{1}$ is the exterior of $P$ in the soid torus $V_{1}$ bounded by $T_{1}$ in $S^{3}$, and $X_{1}$ is the exterior of $K_{1}$. By the above, $T_{1}$ can be isotoped into $X_{0}$. If $T_{1}$ is parallel to $T_{0}$ we have conclusion (a). Otherwise, we have $V\subset V_{1}$, where the core of $V$ is a pattern $P_{1}$ in $V_{1}$. Hence $P=C_{m,n}(P_{1})$. Also, $X_{0}=M_{1}\cup_{T_{1}}X_{1}$, where $M_{1}$ is the exterior of $P_{1}$ in $V_{1}$, so $K_{0}=P_{1}(K_{1})$, giving (b).
@@ 2.4. Reducible and toroidal surgery on satellite knots
In the statement of Theorem 2.9(a), a 00-bridge braid in a solid torus $V$ is a torus knot standardly embedded in $V$. We refer the readers to [22, §2] for details on $1$-bridge braids in $V$, though we note that they are parameterised up to orientation-preserving ambient homeomorphism by triples $(w,b,t)$, where $w\geq 3$ is the number of the strands of the braid, which equals its winding number in $V$, and $21\leq b,t\leq w-$ 2.
@@
(Berge [2], Gabai [21, 22], Gordon [25] and Scharlemann [52]) Let $K=P(K_{0})$ be a satellite knot with pattern $P$ of winding number $w$ and companion $K_{0}$. Set $T=\partial X_{0}$ and suppose that there is a non-meridional slope in $\mathcal{C}(T)$, i.e., $\mathcal{C}(T)\neq\{1/0\}$. Then:normal-:::
@@
    If $K$ is not a cable knot, then $P$ is a $1$-bridge braid with the winding number $w\geq 5$. Further, there is an integer $a$ such that
@@ 3.1. $LO$ and $NLS$ surgery on satellite knots and winding numbers
Let $K=P(K_{0})$ where the winding number of $P$ is zero, and $K(r)=P(r)\cup_{T}X_{0}$ as before. Since $w(P)=0$, $T$ is essential in $K(r)$ by Gabai [19, Corollary 2.5]. By Lemma 2.5, the longitudinal slope of $P(r)$ is $LO$- and $NLS$-detected in $X_{0}$. Then the result follows from Lemma 2.6. ∎
@@
Let $K=P(K_{0})$ where the winding number $w$ of the pattern $P$ is $1$. Write $X(K)=M\cup_{T}X_{0}$ where $X_{0}$ is the exterior of $K_{0}$, $T=\partial X_{0}$ and $\partial M=\partial X(K)\cup T$ as before. Then $K(p/q)$ can be decomposed as $P(p/q)\cup_{T}X_{0}$, where $P(p/q)$ is the result of doing Dehn filling of $M$ with respect to the $p/q$ slope on $\partial X(K)$.
Since $w=1$, $K$ is not a cable knot by Lemma 2.7(3) and $K(p/q)$ is irreducible for all $p/q\in\mathbb{Q}$ by Corollary 2.12. Further, by Theorem 2.13(1), $T$ remains incompressible in $K(p/q)$.
@@
If $K$ is a satellite $L$-space knot then any essential torus in the exterior of $K$ has winding number $2$ or more. ∎
@@ 3.2. $LO$ and $NLS$ surgery on satellite knots from patterns and companions
It is shown in [27] that if $K=P(K_{0})$ is a satellite $L$-space knot, then both $K_{0}$ and $P(U)$ are $L$-space knots, where $U$ is the unknot. Equivalently, if either $K_{0}$ or $P(U)$ is not an $L$-space knot, then $K=P(K_{0})$ is not an $L$-space knot. In this section, we prove the following analogous result (Theorem 3.6) for left-orderability.
@@
First, if $K$ is a cable knot, then by Lemma 2.7(1) the pattern $P$ must be a cabled pattern. Since $\pi_{1}(P(U)(r))$ is left-orderable, then $r$ cannot be the cabling slope. So $K(r)$ is irreducible by Corollary 2.11. Therefore, by [6, Theorem 1.1], it has a left-orderable fundamental group if and only if there is a non-trivial homomorphism of $\pi_{1}(K(r))$ to a left-orderable group. Collapsing the exterior of $K_{0}$ to a solid torus shows that there is a slope-preserving degree one map $(X(K),\partial X(K))\to(X(P(U)),\partial X(P(U)))$, and hence it induces a degree $1$ map $K(r)\to P(U)(r)$. The claim then follows. ∎
@@
Suppose that $w\geq 1$ and $K_{0}(r/w^{2})$ is $LO$. Then $r/w^{2}$ is $LO$-detected in $X_{0}$ by [4, Corollary 8.3]. Also, $\lambda_{P(r)}=r/w^{2}$ by (2.2.6), so if $T=\partial X_{0}$ is incompressible in $K(r)$, then $K(r)$ is $LO$ by Lemma 2.6. On the other hand, if $T$ compresses in $K(r)$, then by Theorem 2.13(1) and (2)(c) we have $K(r)\cong K_{0}(r/w^{2})$ and therefore is $LO$. ∎
@@
Let $M$ be a closed, connected, orientable, irreducible $3$-manifold. For $\ast\in\{NLS,\,CTF\}$, if there exists a nonzero degree map from $M$ to a closed, connected, orientable, irreducible $3$-manifold $N$ that has property $\ast$, does $M$ have property ***?
Since Question 3.9 is known to have a positive answer when *$*=LO$, the $L$-space conjecture predicts positive answers for both $NLS$ and $CTF$, though currently there are very few results in this direction. See [40, 31] for partial answers to Question 3.9 in the $NLS$ case.

### v4

@@ 1. Introduction
We call a closed, connected, orientable, irreducible $3$-manifold $LO$ if its fundamental group is left-orderable  [^1] We adopt the usual convention in the context of the $L$-space conjecture that the trivial group is not left-orderable. See for instance [8, §6.1]. Thus $S^{3}$ is not $LO$., $CTF$ if it admits a co-oriented taut foliation, and $NLS$ if it is not a Heegaard Floer $L$-space. The $L$-space conjecture contends that these three conditions are equivalent ([7], [36]).
@@
Since cable knots are prime, the following corollary is immediate. See Figure 1.
@@
All rational surgeries on a composite knot are $LO$ and $NLS$.
@@
Suppose that $K$ is a satellite $L$-space knot. Then the JSJ graph of $K$ is a rooted interval.
@@
Suppose that $K$ is a satellite knot which admits an irreducible rational surgery that is not $LO$. Then the JSJ graph of $K$ is a rooted interval.
@@
We note that the hypothesis in Theorem 3.13 that the surgery is irreducible is necessary, though not particularly restrictive. A fundamental result of Scharlemann states that surgery on a satellite knot $K$ yields a reducible manifold if and only if $K$ is a cable knot and the surgery slope is the cabling slope ([52]; also see §2.4). As such a surgery has a lens space summand, its fundamental group is not left-orderable. Hence Theorem 3.13 fails without the irreducibility assumption for a knot $K$ if and only if $K$ is a cable of a knot whose JSJ graph is not a rooted interval.
@@
The key result that is used to prove Theorems 3.1 and 3.10 is the fact that the meridional slope of a non-trivial knot $K$ is both $LO$-detected and $NLS$-detected [8] (see Theorem 2.4). (We refer the reader to §2.1 for the definitions of a slope on $\partial X(K)$ being $LO$-, $NLS$- and $CTF$-detected.) The proof of Theorem 3.3 uses the more general fact that any slope on the boundary of an integer homology solid torus at distance 1 from the longitude is $LO$- and $NLS$-detected. In [8, Conjecture 1.6] we conjectured that the latter also holds for $CTF$-detection. In particular,
@@
In [8], we showed that Conjecture 1.1 holds for fibred knots. This allows us to partially extend our results on $LO$ and $NLS$ surgeries to $CTF$ surgeries in the case that the companion knots are fibred. See Theorems 4.1 and 4.2 for the precise statements; these are analogues of Theorems 3.1 and 3.10 respectively. The following analogue of Theorem 3.13 is an immediate consequence of Theorem 4.2.
Theorem 4.3.
Suppose that $K$ is a fibered satellite knot which admits an irreducible rational surgery that is not $CTF$. Then the JSJ graph of $K$ is a rooted interval.
Again, as in Theorem 3.13, the only case where the irreducibility assumption is needed is when $K$ is a cable of a fibered knot whose JSJ graph is not a rooted interval.
Theorem 4.3 implies the following result of Delman and Roberts, which is an immediate consequence of [13, Theorem 6.1].
@@
In addition to what we have mentioned above, it is known that if a satellite knot $K=P(K_{0})$, with pattern $P$ and companion $K_{0}$, is an $L$-space knot, then both $K_{0}$ and $P(U)$ are $L$-space knots [27]. One therefore expects the $LO$ analogue of this to hold also. In this direction, we prove the following.
@@
Suppose that $K$ is a satellite knot with a persistently foliar companion whose associated pattern has winding number $w\geq 1$. Then $K(r)$ is $CTF$ for any $r\in\mathbb{Q}$ which is not the cabling slope if $K$ is a cable knot.
@@
If $K$ is a positive satellite $L$-space knot with pattern of winding number $w$ then $K(r)$ is $CTF$ for each rational $r\in(-\infty,w^{2}]$.
@@
At most two rational surgeries on a non-trivial knot yield integer homology $3$-spheres that are not $LO$, and the associated slopes are contained in either $\{-1,1\}$, $\{-1,-1/2\}$, or $\{1/2,1\}$.
@@ Acknowledgements.
The authors wish to thank the referee for comments which led to an improvement of the exposition and Duncan McCoy and Patricia Sorya for helpful remarks related to Theorem 3.10.
@@ 2.1. Slope detection and gluing results
  • $NLS$-detected (in $M$) if it lies in the closure of the set of slopes whose associated Dehn filling is not an $L$-space ([48, §7.2]). Here the closure is taken inside the set of all rational slopes, which is homeomorphic to $\mathbb{Q}P^{1}$.
@@
It is not known whether a Dehn filling $M(\alpha)$ of a knot manifold $M$ being $CTF$ implies that the slope $\alpha$ is $CTF$-detected, though this is thought to be true (and is true if $M$ is a Seifert fibre space [3]). If so, the condition that $\alpha$ is $CTF$-detected will be strictly weaker than the condition that $M(\alpha)$ be $CTF$: the meridional slope $\mu$ of a non-trivial fibred knot is $CTF$-detected (cf. Theorem 2.4) though the $\mu$-Dehn filling is $S^{3}$ and therefore not $CTF$.
@@
The longitudinal slope of a knot manifold is $\ast$-detected, where * denotes either $NLS$, $CTF$, or $LO$.
@@
Suppose that $W=M_{1}\cup_{f}M_{2}$ where $M_{1},M_{2}$ are knot manifolds and $f:\partial M_{1}\xrightarrow{\;\cong\;}\partial M_{2}$. If $f$ identifies a $\ast$-detected slope on the boundary of $M_{1}$ with a $\ast$-detected slope on the boundary of $M_{2}$, then $W$ has property $\ast$.
@@ 2.2. Conventions on satellite knots
For a coprime pair $p$ and $q$ with $p\geq 0$, define $M(p/q)$ to be the $p\mu+q\lambda$ Dehn filling of $M$ along $\partial X(K)$. Equivalently, $M(p/q)$ is the $p/q$-Dehn surgery on the knot $P$ in $V$. We often denote $M(p/q)$ by $P(p/q)$. Similarly, we use $K(p/q)$ to denote the $p\mu+q\lambda$ Dehn filling of $X(K)$ along $\partial X(K)$. Then
@@
Let $K=P(K_{0})$ be a satellite knot with pattern $P$ and companion $K_{0}$. Fix $r\in\mathbb{Q}$. If $T=\partial X_{0}$ is incompressible in $K(r)$ and $\lambda_{P(r)}$ is *-detected in $X_{0}$, then $K(r)$ has property *, where * is either $LO,NLS$, or $CTF$.
@@ 2.3. Conventions on cable knots
Suppose $K=P(K_{1})$. This gives a decomposition $X(K)=M_{1}\cup_{T_{1}}X_{1}$, where $M_{1}$ is the exterior of $P$ in the solid torus $V_{1}$ bounded by $T_{1}$ in $S^{3}$, and $X_{1}$ is the exterior of $K_{1}$. By the above, $T_{1}$ can be isotoped into $X_{0}$. If $T_{1}$ is parallel to $T_{0}$ we have conclusion (a). Otherwise, we have $V\subset V_{1}$, where the core of $V$ is a pattern $P_{1}$ in $V_{1}$. Hence $P=C_{m,n}(P_{1})$. Also, $X_{0}=M_{1}\cup_{T_{1}}X_{1}$, where $M_{1}$ is the exterior of $P_{1}$ in $V_{1}$, so $K_{0}=P_{1}(K_{1})$, giving (b).
@@ 2.4. Reducible and toroidal surgery on satellite knots
In the statement of Theorem 2.9(a), a 00-bridge braid in a solid torus $V$ is a torus knot standardly embedded in $V$, in other words, a cable pattern $C_{m,n}$. We refer the readers to [22, §2] for details on $1$-bridge braids in $V$, though we note that they are parameterised up to orientation-preserving ambient homeomorphism by triples $(w,b,t)$, where $w\geq 3$ is the number of the strands of the braid, which equals its winding number in $V$, and $21\leq b,t\leq w-$ 2.
@@
(Berge [2], Gabai [21, 22], Gordon [25] and Scharlemann [52]) Let $K=P(K_{0})$ be a satellite knot with pattern $P$ of winding number $w$ and companion $K_{0}$. Set $T=\partial X_{0}$ and suppose that there is a non-meridional slope in $\mathcal{C}(T)$, i.e., $\mathcal{C}(T)\neq\{1/0\}$. Then::::
@@
    If $K$ is not a cable knot, then $P$ is a $1$-bridge braid with winding number $w\geq 5$. Further, there is an integer $a$ such that
@@
Lemma 2.14.
Let $K=P(K_{0})$ where $P$ has winding number $w$ and $X_{0}$ is the exterior of $K_{0}$. Then $K(w^{2})$ is irreducible and $\partial X_{0}$ is incompressible in $K(w^{2})$.
Proof.
The first conclusion follows from the second by Corollary 2.11 and Theorem 2.13. Also by that theorem, if $\partial X_{0}$ compresses in $K(r)$ and $r$ is integral then either $P=C_{m,n}$ and $r=mn$ or $mn\pm 1$, or $P$ is a 1-bridge braid and $P(r)$ is a solid torus.
In the first case $r\neq m^{2}$. In the second case, by [22] $r=a$ or $a+1$ where $a=tw+b$ and $21\leq b,t\leq w-$ 2. Then $a+1\leq w(w-2)+(w-2)+1=w^{2}-w-1<w^{2}$. ∎
@@ 3.1. $LO$ and $NLS$ surgery on satellite knots and winding numbers
Let $K=P(K_{0})$ where the winding number of $P$ is zero, and $K(r)=P(r)\cup_{T}X_{0}$ as before. Since $w(P)=0$, $T$ is essential in $K(r)$ by Gabai [19, Corollary 2.5]. Since $r$ is not the cabling slope, by Theorem 2.9, $P(r)$ is irreducible. Hence by Proposition 2.2, the longitudinal slope of $P(r)$ is $LO$- and $NLS$-detected in $P(r)$. By Lemma 2.5, the longitudinal slope of $P(r)$ is also $LO$- and $NLS$-detected in $X_{0}$. Then the result follows from Lemma 2.6. ∎
@@
Let $K=P(K_{0})$ where the winding number $w$ of the pattern $P$ is $1$. Write $X(K)=M\cup_{T}X_{0}$ where $X_{0}$ is the exterior of $K_{0}$, $T=\partial X_{0}$ and $\partial M=\partial X(K)\cup T$ as before. Then $K(p/q)$ can be decomposed as $P(p/q)\cup_{T}X_{0}$.
Since $w=1$, $K(p/q)$ is irreducible for all $p/q\in\mathbb{Q}$ by Corollary 2.12. Further, by Theorem 2.13(1), $T$ remains incompressible in $K(p/q)$.
@@
If $K$ is a satellite $L$-space knot then any essential torus in the exterior of $K$ has winding number $2$ or more.
@@ 3.2. $LO$ and $NLS$ surgery on satellite knots from patterns and companions
It is shown in [27] that if $K=P(K_{0})$ is a satellite $L$-space knot, then both $K_{0}$ and $P(U)$ are $L$-space knots, where $U$ is the unknot. Equivalently, if either $K_{0}$ or $P(U)$ is not an $L$-space knot, then $K=P(K_{0})$ is not an $L$-space knot. In this section, we prove the following analogous result for left-orderability.
@@
First, if $K$ is a cable knot, then by Lemma 2.7(1) the pattern $P$ must be a cabled pattern. Since $\pi_{1}(P(U)(r))$ is left-orderable, then $r$ cannot be the cabling slope. So $K(r)$ is irreducible by Corollary 2.11. Therefore, by [6, Theorem 1.1], it has a left-orderable fundamental group if and only if there is a non-trivial homomorphism of $\pi_{1}(K(r))$ to a left-orderable group.
Let $F_{0}$ be the Seifert surface of $K_{0}$. We can collapse the exterior $X(K_{0})$ of $K_{0}$ to a solid torus by collapsing a regular neighborhood of $\partial X(K_{0})\cup F_{0}$ to a solid torus with an interior point removed and mapping the remaining part of $X(K_{0})$ to the removed point. This shows that there is a slope-preserving degree one map $(X(K),\partial X(K))\to(X(P(U)),\partial X(P(U)))$, and hence it induces a degree $1$ map $K(r)\to P(U)(r)$. The claim then follows. ∎
@@
Suppose that $w\geq 1$ and $K_{0}(r/w^{2})$ is $LO$. Then $r/w^{2}$ is $LO$-detected in $X_{0}$ by [4, Corollary 8.3]. Also, $\lambda_{P(r)}=r/w^{2}$ by (2.2.6), so if $T=\partial X_{0}$ is incompressible in $K(r)$, then $K(r)$ is $LO$ by Lemma 2.6. On the other hand, since $r$ is not the cabling slope, the case in Theorem 2.13(2)(b) doesn’t occur. Hence, if $T$ compresses in $K(r)$, then by Theorem 2.13(1) and (2)(c) we have $K(r)\cong K_{0}(r/w^{2})$ and therefore is $LO$. ∎
@@
Let $M$ be a closed, connected, orientable, irreducible $3$-manifold. For $\ast\in\{NLS,\,CTF\}$, if there exists a nonzero degree map from $M$ to a closed, connected, orientable, irreducible $3$-manifold $N$ that has property $\ast$, does $M$ have property *?
Since Question 3.9 is known to have a positive answer when $*=LO$ [6, Theorem 3.7], the $L$-space conjecture predicts positive answers for both $NLS$ and $CTF$, though currently there are very few results in this direction. See [40, 31] for partial answers to Question 3.9 in the $NLS$ case.